OpenStudy (anonymous):

can someone please help me to find the interquartile range of 6,10,12,4,3,7,15,20

jimthompson5910 (jim_thompson5910):

first sort the numbers

jimthompson5910 (jim_thompson5910):

from least to greatest

jimthompson5910 (jim_thompson5910):

what do you get when you do this

OpenStudy (anonymous):

so it would be 3,4,6,7,10,12,15,20

jimthompson5910 (jim_thompson5910):

good it's 3, 4, 6, 7, 10, 12, 15, 20

jimthompson5910 (jim_thompson5910):

now start at each endpoint and cross out the pair of outermost terms so start with 3, 4, 6, 7, 10, 12, 15, 20 and cross out 3 and 20 to get 4, 6, 7, 10, 12, 15

jimthompson5910 (jim_thompson5910):

repeat the last step and cross out 4 and 15 6, 7, 10, 12

jimthompson5910 (jim_thompson5910):

cross out the 6 and 12 to get 7, 10

jimthompson5910 (jim_thompson5910):

the middle most value will be between 7 and 10 so the median is (7+10)/2 = 17/2 = 8.5

jimthompson5910 (jim_thompson5910):

now we have this data set 3, 4, 6, 7, 10, 12, 15, 20 which has a median of 8.5 break the data set into two halves like this 3, 4, 6, 7 10, 12, 15, 20 and find the median of the lower and upper halves